Fermi National Laboratory
 

Operation Summary for October 26, 2011 - October 28, 2011

 

Wednesday October 26

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the MiniBooNE and NuMI experiments taking beam.

During the day shift, RF personnel tried to bring a MI RF station (MIRF17) back online at 8:26 AM, but it continued to trip off; RF personnel investigated. Antiproton Source (Pbar) personnel began a study around 8:30 AM. At 3:17 PM, the NuMI beam permit tripped off due to an incomplete RF gap, but no RF stations had tripped off; it reset, but then tripped off again. Operators investigated and found MIRF17 was the cause; RF personnel investigated.

During the evening shift, the NuMI beam permit tripped off at 4:59 PM when the output of a quadrupole power supply (E:QF107) drifted low; everything reset. Operations established beam to the Meson FTBF experiment T-1015 at 5:01 PM. Pbar studies ended at 5:35 PM. Operations disabled beam to T-1015 at 10:02 PM. Operations reported a short downtime due to a Linac RF station (LRF1).

Thursday October 27

The midnight shift began with Operations monitoring beam to the NuMI and MiniBooNE experiments.

During the midnight shift, Operations disabled beam to MiniBooNE at 5 AM in preparation for an access later in the morning. At 7:45 AM, Neutron Therapy Facility (NTF) personnel began preparing to treat two patients later in the morning.

During the day shift, Operators prepared the MiniBooNE MI-12 enclosure at 8:45 AM for an access. Operations established beam to T-1015 at 10:30 AM. The MI-12 access ended and operators tried to send beam to MiniBooNE, but a problem with a quadrupole (E:Q874) power supply delayed things. Operations established beam to MiniBooNE at 10:49 AM. Operations turned off the Pbar rings, Pretarget, and Prevault power supplies at 3:20 PM.

During the evening shift, a Booster RF station (BRF4) tripped off for the first of two times in 19 minutes; it reset, but a RF technician investigated. Operations disabled beam to T-1015 at 10:05 PM.

Friday October 28

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE experiments.

During the midnight shift, operators turned on the Pbar power supplies at 3:41 AM in preparation for a study. At 3:49 AM, the I- Source voltage tripped off along with two quadrupole power supplies (QPS105 and 108); operators turned everything back on. The NuMI beam permit tripped off at 3:58 AM due to the output of E:QPS114 drifting low; everything reset. Operators reported conducting a 750 KeV line tune at 4:05 AM. Operators began tuning up Pbar for reverse proton injection into the Accumulator. A Pbar person began a Mu2e study at 4:51 AM. At 7:50 AM, NTF personnel began preparing to treat two patients later in the morning. Operators prepared the Meson enclosure M05 for an access at 7:59 AM.

Operations Coordinator (OpsCo) Plans for Today and the Weekend

The plans for today include Pbar studies and a MI-60 access. The plans for today and the weekend are to keep delivering protons to the experiments. NuMI personnel reported that on Wednesday, 10/26/11, they set a record for the most recorded protons on target (POT) for one day, 1.4391E18, and then on Thursday, 10/27/11, they recorded the second highest POT for one day.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/28/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory